# Ensembling Prioritized Hybrid Policies for Multi-agent Pathfinding

### v1

@@ Abstract
Multi-Agent Reinforcement Learning (MARL) based Multi-Agent Path Finding (MAPF) has recently gained attention due to its efficiency and scalability. Several MARL-MAPF methods choose to use communication to enrich the information one agent can perceive. However, existing works still struggle in structured environments with high obstacle density and a high number of agents. To further improve the performance of the communication-based MARL-MAPF solvers, we propose a new method, Ensembling Prioritized Hybrid Policies (EPH). We first propose a selective communication block to gather richer information for better agent coordination within multi-agent environments and train the model with a Q-learning-based algorithm. We further introduce three advanced inference strategies aimed at bolstering performance during the execution phase. First, we hybridize the neural policy with single-agent expert guidance for navigating conflict-free zones. Secondly, we propose Q value-based methods for prioritized resolution of conflicts as well as deadlock situations. Finally, we introduce a robust ensemble method that can efficiently collect the best out of multiple possible solutions. We empirically evaluate EPH in complex multi-agent environments and demonstrate competitive performance against state-of-the-art neural methods for MAPF.
@@ I Introduction
Multi-Agent Reinforcement Learning (MARL) based approaches [11, 12, 13, 14] offer another way to solve the MAPF problem. Instead of treating MAPF as a centralized problem, they solve MAPF by regarding it as a sequential decision-making problem. One of the seminal MARL-based approaches to MAPF is PRIMAL[11]. In PRIMAL, the authors propose to learn a fully decentralized policy with partial agent-wise observation; the model is trained with A3C [15] and imitation learning (IL), However, PRIMAL struggles to scale to high-obstacle-density environments with many agents. Also, PRIMAL assumes that agents move sequentially in each time step, which differs from real-world scenarios where agents move simultaneously in each time step. The improved version of PRIMAL, PRIMAL2 [12], also trains with A3C and IL. PRIMAL2 assumes the disappear at target [1] scenario. This assumption makes the problem easier because it means lower obstacle density when agents reach their goals and disappear, whereas in the stay at target [1] setting, agents constitute obstacles since they do not disappear after reaching their goals. Thus, the problem is much harder, especially when many agents exist. We adopt stay at target in this work.
@@
Contributions. In this paper, we propose EPH (Ensembling Prioritized Hybrid Policies), a Q-learning-based MARL-MAPF solver with communication. Our communication scheme entails enhanced selective communication block with improvements inspired by the latest Transformer variant to enable richer information extraction. Additionally, we introduce several strategies to be used during the inference phase to further bolster performance. Firstly, we propose Q value-based priority decisions, in which the Prioritized Conflict Resolution decides the priority of agents involved in conflicts, and Advanced Escape Policy is responsible for breaking deadlocks. Secondly, we use hybrid expert guidance for agents that have no other live agents nearby. Lastly, we utilize ensembling to sample the best solutions from multiple solvers that run in parallel to leverage the different strengths of each strategy proposed. By improving communication capability and utilizing proposed strategies to help inference, we achieve competitive performance against state-of-the-art neural MARL-MAPF solvers.
@@ III-B Environment Settings
Following the conventions of MAPF, we use a 2D grid world to represent the environment. Each cell in the $n\times n$ grid world can be either an empty cell or an obstacle that stops agents from passing [^1] We consider value $1$ on map $M$ as an obstacle and 00 as available.. Each agent occupies an empty cell, and the task for each agent is moving from its starting cell to the goal cell without collisions. At the beginning of each episode, $m$ starting cells and $m$ goal cells are randomly selected among the empty cells for $m$ agents, and we make sure that there is no overlap among $2m$ selected cells. At each timestep, the agents choose to move either up, down, left, right, or stay still. We consider four types of conflicts: swapping conflict [1], vertex conflict [1], conflict with static obstacles, and out-of-bound conflict. These four types of conflicts are not allowed during pathfinding.
@@ IV EPH: Ensembling Prioritized Hybrid Policies
We propose EPH and describe it in two sections, each with several subsections. Section IV-A describes how we formulate MAPF problem as a Q-Learning based MARL problem with communication and how we train such a model. Section IV-B gives introduction about the advanced inference strategies that we propose to improve performance during execution. Figure 1 shows the overall architecture of EPH.
@@ IV-A1 Graph Convolution based Communication
We show our communication architecture in Figure 1. Our communication structure is adapted from selective communication block in DCC [13]. Compared with ordinary communication, selective communication can effectively reduce communication time compared with non-selective counterparts, it can also selectively gather more informative and important hidden information from other agents within FOV, which is shown to be beneficial for pathfinding [13]. However, in the previous work, the selective graph convolution communication block did not include position-wise processing or normalization, and it is shown that including them helps stabilize training and gives better scalability [29]. Inspired by GTrXL-styled transformer in [29], we re-design the communication block to enable better coordination among agents by including extra multi-layer perception layer, GRU block, and layer normalization, which is shown to be beneficial in many works across different fields [30, 17, 29]. The generation method of communication mask for selective communication, as well as the request-reply communication scheme that consists of two cascaded communication blocks, are kept the same as shown in DCC to reduce communication overhead and enrich the information gathered.
@@ IV-B Advanced Inferences Strategies for EPH
We propose three three inference techniques to further bolster pathfinding performance during inference. In Section IV-B1, we hybridize the policy with single-agent optimal $A^{*}$ paths to give guidance to agents that do not have any other live agents nearby. In Section IV-B2, we introduce two Q value-based priority decision-making strategies, namely Prioritized Conflict Resolution for efficient conflict resolution and Advanced Escape Policy for priority-based avoidance of deadlocks. Finally, Section IV-B3 introduces the ensembling, which runs multiple policies in parallel and samples the best possible solutions in the solution space.
@@ IV-B1 Hybrid Expert Guidance
When agent communication is sparse, such as in scenarios where no other agents are located within the agent’s field of view, it is beneficial to incorporate low-cost single-agent expert path to guide decision-making, thus hybridizing (EPH stands for Hybrid). We hereby define a live agent as an agent that is currently off its goal, i.e., $v_{i}^{t}\neq v_{i}^{g}$. We propose leveraging the optimal $A^{*}$ path as expert guidance during inference for an agent that has no other live agents within a square centered on the agent itself; the length of the square is $12\rho+$ 1 where $\rho$ is the visibility radius for live agents. We adapt the efficient $A^{*}$, a well-established low-cost algorithm for efficient single-agent pathfinding, to multi-agent settings with the following formulation:
@@ IV-B2 Value-based Priority Decisions
We further introduce two techniques for resolving conflict situations and deadlock scenarios based on priorities (EPH stands for Prioritized) provided by the Q values obtained by the trained neural model. This is inspired by priority-based planning methods [34, 35, 36] and, in particular, Priority-Based Search (PBS) [37], which first decides the priority for each agent and then plans individual paths for each agent in the order of priority. In PBS, the agent with lower priority is not allowed to collide with the agent with higher priority. In our Q value-based priority decisions, an agent with a lower Q value will be assigned a lower priority and will have to give way to another agent with a higher priority.
@@
Gao et al. [27] first proposed an escape policy to avoid deadlock situations, which is a common case in structured environments. A deadlock can occur when an agent is off its goal and oscillates back and forth between two adjacent cells (Algorithm 1, lines 9-10). However, in the original escape policy in [27], agents choose to break deadlocks just by taking random actions when deadlocks are detected, which can be suboptimal and lead to unnecessary randomness. In our Advanced Escape Policy, we use a two-stage approach to break the deadlocks. Firstly, we assign a priority to each agent by sorting the Q values in descending order. Then, for each agent in such order, we utilize $A^{*}_{\tau}$ mentioned in Section IV-B1 to obtain the action. If $A^{*}_{\tau}$ fails to find a path because of other agents involved in the deadlock, we choose the action from the highest valid Q value, where invalid Q values (corresponding to actions that lead to a collision) are masked, i.e. set to $\infty$- ∞. Notably, we also update a copy of the map such that the current next action is masked, ensuring there will be no conflict. Algorithm 1 illustrates our proposed Advanced Escape Policy.
@@ 17                         $a^{t}_{i}\leftarrow\text{arg}\max{q_{i}^{t}}$
(a) Success Rate
(b) Average Episode Length
Figure 4: Comparative Analysis of Success Rate and Average Episode Length on random maps.
@@ V-A Training and Testing Settings
We use D3QN with a prioritized experience replay buffer [38] for training EPH. Curriculum learning [39], which is shown to be effective in previous works [14, 18], is also adopted. The curriculum learning starts from one agent in $10\times 10$ map, and ends on $40\times 40$ map with 16 agents. Every time the success rate of the model under training reaches 0.9 for the current training environment, we increase the difficulty of the map. Thus, experience from more complex environments will be added to the buffer, and the model learns from more and more complex scenarios gradually. The obstacle density of maps in training is sampled from a triangular distribution between 0 and 0.5 with a peak of 0.33. The FOV size in our work is $9\times 9$ to align with previous works [14, 13]. The maximum episode length for training episodes is 256. We save model checkpoints every 1000 training steps and take the best one as the final model based on performance on a random-map validation set. It takes 20 hours to perform 150k training steps on a server with a single NVIDIA RTX A6000 and AMD EPYC 7542 (128) @2.9GHz. We also retrain DCC with the same setting for fair comparison. We value open reproducibility and make our code publicly available [^2] https://github.com/ai4co/eph-mapf.
@@ V-B Results on Random Maps
Figure 3(a) shows the success rate of EPH compared with baselines. On $40\times 40$ map, EPH beats all the baselines in terms of success rate, reaching $100\%$ success rate for all the cases; on the $80\times 80$ map, EPH outperforms both DHC and DCC in all cases, while falling behind SCRIMP by $1\%$ on 32 and 64 agents scenarios. Figure 3(b) illustrates the average episode length, which indicates the solution quality, of all four models. EPH outperforms both DHC and DCC in terms of average episode length in all cases. When compared against SCRIMP, EPH can outperform SCRIMP on larger $80\times 80$ map. On $40\times 40$ map, EPH can achieve a shorter average episode length than SCRIMP except in the 64-agent case. It should be noted that SCRIMP uses global communication, which could give the model more information than we could. However, global communication is burdensome and inefficient in real life; also, global communication may be limited by geographical constraints. By contrast, EPH, which achieves better performance than baselines in most cases, only uses improved local selective communication. EPH equipped with local selective communication is information-efficient, less burdensome, and practical in real-life scenarios.
@@ V-C Results on Structured Maps
We additionally test on structured maps that were never seen during training to showcase the performance of baselines and EPH in terms of scalability and generalization to unseen scenarios. For structured maps, apart from the abovementioned baselines, we further include three additional heuristic solvers: CBS [7], wPBS [10] and $*{}^{*}$ [41]; as well as two neural solvers: PRIMAL [11] and the recent SACHA [18] as baselines. SACHA is a neural MAPF solver trained with Soft Actor-Critic and employs optional global communication blocks. We choose to compare with SACHA equipped with global communication block for fair comparison. For the heuristics solvers, the runtime limit is 120 seconds for CBS and wPBS, and 20 seconds for $*{}^{*}$, same as the settings used in SACHA.

### v2

@@ Abstract
Multi-Agent Reinforcement Learning (MARL) based Multi-Agent Path Finding (MAPF) has recently gained attention due to its efficiency and scalability. Several MARL-MAPF methods choose to use communication to enrich the information one agent can perceive. However, existing works still struggle in structured environments with high obstacle density and a high number of agents. To further improve the performance of the communication-based MARL-MAPF solvers, we propose a new method, Ensembling Prioritized Hybrid Policies (EPH). We first propose a selective communication block to gather richer information for better agent coordination within multi-agent environments and train the model with a Q-learning-based algorithm. We further introduce three advanced inference strategies aimed at bolstering performance during the execution phase. First, we hybridize the neural policy with single-agent expert guidance for navigating conflict-free zones. Secondly, we propose Q value-based methods for prioritized resolution of conflicts as well as deadlock situations. Finally, we introduce a robust ensemble method that can efficiently collect the best out of multiple possible solutions. We empirically evaluate EPH in complex multi-agent environments and demonstrate competitive performance against state-of-the-art neural methods for MAPF. We open-source our code at https://github.com/ai4co/eph-mapf.
@@ I Introduction
Multi-Agent Reinforcement Learning (MARL) based approaches [11, 12, 13, 14] offer another way to solve the MAPF problem. Instead of treating MAPF as a centralized problem, they solve MAPF by regarding it as a sequential decision-making problem. One of the seminal MARL-based approaches to MAPF is PRIMAL[11]. In PRIMAL, the authors propose to learn a fully decentralized policy with partial agent-wise observation; the model is trained with MARL and imitation learning (IL), However, PRIMAL struggles to scale to high-obstacle-density environments with many agents. Also, PRIMAL assumes that agents move sequentially in each time step, which differs from real-world scenarios where agents move simultaneously in each time step. The improved version of PRIMAL, PRIMAL2 [12], also trains with MARL and IL. PRIMAL2 assumes the disappear at target [1] scenario. This assumption makes the problem easier because it means lower obstacle density when agents reach their goals and disappear, whereas in the stay at target [1] setting, agents constitute obstacles since they do not disappear after reaching their goals. Thus, the problem is much harder, especially when many agents exist. We adopt stay at target in this work.
@@
Contributions. In this paper, we propose EPH (Ensembling Prioritized Hybrid Policies), a Q-learning-based MARL-MAPF solver with communication. Our communication scheme entails an enhanced selective communication block with improvements inspired by the latest Transformer variant to enable richer information extraction. Additionally, we introduce several strategies to be used during the inference phase to further bolster performance. Firstly, we propose Q value-based priority decisions, in which the Prioritized Conflict Resolution decides the priority of agents involved in conflicts, and Advanced Escape Policy is responsible for breaking deadlocks. Secondly, we use hybrid expert guidance for agents that have no other live agents nearby. Lastly, we utilize ensembling to sample the best solutions from multiple solvers that run in parallel to leverage the different strengths of each strategy proposed. By improving communication capability and utilizing proposed strategies to help inference, we achieve competitive performance against state-of-the-art neural MARL-MAPF solvers.
@@ III-B Environment Settings
Following the conventions of MAPF, we use a 2D grid world to represent the environment. Each cell in the $n\times n$ grid world can be either an empty cell or an obstacle that stops agents from passing [^1] We consider value $1$ on map $M$ as an obstacle and 00 as available.. Each agent occupies an empty cell, and the task for each agent is moving from its starting cell to the goal cell without collisions. At the beginning of each episode, $m$ starting cells and $m$ goal cells are randomly selected among the empty cells for $m$ agents, and we make sure that there is no overlap among $2m$ selected cells. At each timestep, the agents choose to move either up, down, left, right, or stay still. We consider four types of conflicts: edge conflict [1], vertex conflict [1], conflict with static obstacles, and out-of-bound conflict. These four types of conflicts are not allowed during pathfinding.
@@ IV EPH: Ensembling Prioritized Hybrid Policies
We propose EPH and describe it in two sections, each with several subsections. Section IV-A describes how we formulate MAPF problem as a Q-Learning based MARL problem with communication and how we train such a model. Section IV-B gives an introduction to the advanced inference strategies that we propose to improve performance during execution. Figure 1 shows the overall architecture of EPH.
@@ IV-A1 Graph Convolution based Communication
We show our communication architecture in Figure 1. Our communication structure is adapted from selective communication block in DCC [13]. Compared with ordinary communication, selective communication can effectively reduce communication time compared with non-selective counterparts, it can also selectively gather more informative and important hidden information from other agents within FOV, which is shown to be beneficial for pathfinding [13]. However, in the previous work, the selective graph convolution communication block did not include position-wise processing or normalization, and it is shown that including them helps stabilize training and gives better scalability [28]. Inspired by GTrXL-styled transformer in [28], we re-design the communication block to enable better coordination among agents by including an extra multi-layer perception layer, GRU block, and layer normalization, which is shown to be beneficial in many works across different fields [29, 16, 28]. The generation method of communication mask for selective communication, as well as the request-reply communication scheme that consists of two cascaded communication blocks, are kept the same as shown in DCC to reduce communication overhead and enrich the information gathered.
@@ IV-B Advanced Inferences Strategies for EPH
We propose three inference techniques to further bolster pathfinding performance during inference. In Section IV-B1, we hybridize the policy with single-agent optimal $A^{*}$ paths to give guidance to agents that do not have any other live agents nearby. In Section IV-B2, we introduce two Q value-based priority decision-making strategies, namely Prioritized Conflict Resolution for efficient conflict resolution and Advanced Escape Policy for priority-based avoidance of deadlocks. Finally, Section IV-B3 introduces the ensembling, which runs multiple policies in parallel and samples the best possible solutions in the solution space.
@@ IV-B1 Hybrid Expert Guidance
When agent communication is sparse, such as in scenarios where no other agents are located within the agent’s field of view, it is beneficial to incorporate low-cost single-agent expert path to guide decision-making, thus hybridizing (EPH stands for Hybrid). We hereby define a live agent as an agent that is currently off its goal, i.e., $v_{i}^{t}\neq v_{i}^{g}$. We propose leveraging the optimal $A^{*}$ path as expert guidance during inference for an agent that has no other live agents within a square centered on the agent itself; the length of the square is $12\rho+$ 1 where $\rho$ is the visibility radius for live agents. We adapt the efficient $A^{*}$, a well-established low-cost algorithm for efficient single-agent pathfinding, to multi-agent settings as follows:
@@ IV-B2 Value-based Priority Decisions
We further introduce two techniques for resolving conflict situations and deadlock scenarios based on priorities (EPH stands for Prioritized) provided by the Q values obtained by the trained neural model. This is inspired by priority-based planning methods [33, 34, 35] and, in particular, Priority-Based Search (PBS) [36], which first decides the priority for each agent and then plans individual paths for each agent in the order of priority. In PBS, the agent with lower priority is not allowed to collide with the agent with higher priority. In our Q value-based priority decisions, an agent with a lower Q value will be assigned a lower priority and will have to give way to another agent with a higher priority. Intuitively, higher Q-values correspond to higher expected rewards.
@@
[…]
@@ 22                         $a^{t}_{i}\leftarrow\text{arg}\max{q_{i}^{t}}$
Figure 4: Comparative Analysis of Success Rate (left) and Average Episode Length (right) on random maps.
@@ V-A Training and Testing Settings
We use D3QN with a prioritized experience replay buffer [37] for training EPH. Curriculum learning [38], which is shown to be effective in previous works [14, 17], is also adopted. The curriculum learning starts from one agent in $10\times 10$ map, and ends on $40\times 40$ map with 16 agents. Every time the success rate of the model under training reaches 0.9 for the current training environment, we increase the difficulty of the map. Thus, experience from more complex environments will be added to the buffer, and the model learns from more and more complex scenarios gradually. The obstacle density of maps in training is sampled from a triangular distribution between 0 and 0.5 with a peak of 0.33. The FOV size in our work is $9\times 9$ to align with previous works [14, 13]. The maximum episode length for training episodes is 256. We save model checkpoints every 1000 training steps and take the best one as the final model based on performance on a random-map validation set. It takes 20 hours to perform 150k training steps on a server with a single NVIDIA RTX A6000 and AMD EPYC 7542 (128) @2.9GHz. We also retrain DCC with the same setting for fair comparison.
@@ V-B Results on Random Maps
Figure 4 (left) shows the success rate of EPH compared with baselines. On $40\times 40$ map, EPH beats all the baselines in terms of success rate, reaching $100\%$ success rate for all the cases; on the $80\times 80$ map, EPH outperforms both DHC and DCC in all cases, while falling behind SCRIMP by $1\%$ on 32 and 64 agents scenarios. Figure 4 (right) illustrates the average episode length, which indicates the solution quality, of all four models. EPH outperforms both DHC and DCC in terms of average episode length in all cases. When compared against SCRIMP, EPH can outperform SCRIMP on larger $80\times 80$ map. On $40\times 40$ map, EPH can achieve a shorter average episode length than SCRIMP except in the 64-agent case. It should be noted that SCRIMP uses global communication, which could give the model more information than we could. However, global communication is burdensome and inefficient in real life; also, global communication may be limited by geographical constraints. By contrast, EPH, which achieves better performance than baselines in most cases, only uses improved local selective communication. EPH equipped with local selective communication is information-efficient, less burdensome, and practical in real-life scenarios.
@@ V-C Results on Structured Maps
We additionally test on structured maps that were never seen during training to showcase the performance of baselines and EPH in terms of scalability and generalization to unseen scenarios. For structured maps, apart from the abovementioned baselines, we further include three additional heuristic solvers: CBS [7], wPBS [10] and ODrM [40]; as well as two neural solvers: PRIMAL [11] and the recent SACHA [17] as baselines. SACHA is a neural MAPF solver trained with Soft Actor-Critic and employs optional global communication blocks. We choose to compare with SACHA equipped with global communication block for fair comparison. For the heuristics solvers, the runtime limit is 120 seconds for CBS and wPBS, and 20 seconds for ODrM, same as the settings used in SACHA. EPH takes a slightly higher runtime compared to DCC. At the same time, conflict checks and running A have a negligible effect on runtime [^3] We note that we did not parallelize ensembling in the implementation; parallelizing ensembling would significantly speed up the algorithm. We leave this implementation for future work..